# 3DGS-LM: Faster Gaussian-Splatting Optimization with Levenberg-Marquardt

### v1

@@ 4.2 Ablations
3DGS [20] (L1, SSIM) 0.862 27.23 0.108 1573
3DGS + Ours (L1, SSIM) 0.863 27.29 0.110 1175
@@
Table 2: Ablation of objective. We compare using the L1/SSIM losses against the L2 loss. For both, 3DGS [20] optimized with ADAM and combined with ours, we achieve better results with the L1/SSIM objective. In both cases, our method accelerates the convergence. We report results on the garden scene from the MipNeRF360 [4] dataset starting from the same initialization.
We utilize the same loss functions in our LM optimizer as in the original 3DGS implementation, namely the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses (see Eq. 1). Since LM energy terms are defined as a sum of squares, we adopt the square root formulation of these loss functions to arrive at an identical objective (see Eq. 2). We compare this choice against fitting the Gaussians with only an $\mathcal{L}_{2}$ loss, that does not require taking a square root. Concretely, we compare the achieved quality and runtime of LM against ADAM for both the $\mathcal{L}_{2}$ loss and the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses. As can be seen in Tab. 2, we achieve faster convergence and similar quality in both cases. However, the achieved quality is inferior for both LM and ADAM when only using the $\mathcal{L}_{2}$ loss. This highlights the importance of the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ loss functions and why we adopt them in our method as well.
How many images per batch are necessary?
@@
Table 3: Ablation of batch-size. We vary the number of iterations (#I) and number of images in a batch (#B). Reducing #B for our method reduces runtime and consumed memory, while only slightly impacting quality. This demonstrates that image subsampling Sec. 3.3.2 is compatible with LM in our task. Increasing #I in 3DGS [20] with ADAM is slower and achieves worse quality than Ours using LM. We report average results on the NeRF-Synthetic [28] dataset starting from the same initialization.
The key hyperparameters in our model are the number of images in a batch and how many batches to choose for every LM iteration (Sec. 3.3.2). This directly controls how much GPU memory our optimizer consumes and how much time we need per iteration. We compare different number of images in Tab. 3 (top) on the NeRF-Synthetic [28] dataset. We choose this dataset because it is possible to construct the cache for all 100 training images in 33 GB of GPU memory and optimize the scene with one batch of all images. We compare this against our proposed subsampling scheme for different number of images in a single batch. Decreasing the number of images in a batch results in only slightly worse quality, but also yields faster convergence and reduces GPU memory consumption linearly down to 15GB for 40 images. This demonstrates that subsampling images does not negatively impact the convergence of the LM optimizer in our task. This motivates our usage of subsampling on the real-world datasets in our main results.
Are we better than full-batch ADAM? One reason why LM requires drastically fewer iterations to converge than ADAM is the larger batch size (number of training images per iteration). Concretely, we require only 5 additional LM iterations after the initialization, whereas ADAM runs for another 10K iterations to converge. We can similarly increase the batch size in ADAM to reduce the number of iterations. However, as can be seen in Tab. 3 (bottom), the achieved quality is worse for the same batch size and number of iterations. When running for more iterations, ADAM eventually converges to similar quality, but needs more time. This highlights the efficiency of our optimizer: since we solve the normal equations in Eq. 2, one LM iteration makes a higher quality update step than ADAM which only uses the gradient direction.
4.3 Runtime Analysis
@@
Figure 6: Runtime Analysis. One iteration of our LM optimizer is dominated by solving PCG and building the cache. Measured on the “garden” scene from Mip-NeRF360 [4] after densification.
We analyze the runtime of our LM optimizer across multiple iterations in Fig. 6. The runtime is dominated by solving Eq. 3 with PCG and building the cache (Sec. 3.3). Sorting the cache, rendering all selected images, and the line search (Eq. 4) are comparably faster. When running PCG, we execute the applyJ and applyJT kernels up to 8 times using the per-pixel-per-splat parallelization pattern (Algorithm 1). In contrast, we execute the buildCache kernel once and parallelize per-pixel, which is only marginally faster. This demonstrates the speed advantage obtained by using our proposed caching structure. We also provide a detailed profiling analysis of the kernels in the supplement.
4.4 Limitations
By replacing ADAM with our LM scheme, we accelerate the convergence speed by 30% on average for all datasets and baselines. However, some drawbacks remain.
@@ 4.4 Limitations
First, our approach has a higher GPU memory footprint than baselines, due to our gradient cache (Sec. 3.3). Depending on the number of images and their resolution, this can make it hard to run our method on smaller GPUs. Following Mallick et al. [27], one could reduce the cache size by storing the gradients $\frac{\partial p}{\partial s}$ only for every 32nd splat along a ray and re-doing the $\alpha$-blending in these local windows.
Second, our two-stage approach relies on ADAM to first complete the densification. The original 3DGS [20] densifies Gaussians up to 140 times, which is not easily transferable to the granularity of only 5 LM iterations. Instead, one could explore recent improvements in densification and integrate them into our method [22, 5].
@@ 5 Conclusion
We have presented 3DGS-LM, a method that accelerates the reconstruction of 3D Gaussian-Splatting [20] by replacing the ADAM optimizer with a tailored Levenberg-Marquardt (LM) (Sec. 3.2). We show that with our data parallization scheme we can efficiently solve the normal equations with PCG in custom CUDA kernels (Sec. 3.3). Employed in a two-stage approach (Sec. 3.4), this leads to a 30% runtime acceleration compared to baselines. We further demonstrate that our approach is agnostic to other methods  [11, 41, 27], which further improves the optimization runtime. Overall, we believe that the ability of faster 3DGS reconstructions with our method will open up further research avenues and make existing 3DGS more practical across a wide range of real-world applications.
@@ Appendix A More Details about CUDA Kernel Design
We introduce the necessary CUDA kernels to calculate the PCG algorithm in Sec. 3.3.1. In this section, we provide additional implementation details.
The necessary computations for the buildCache and applyJT steps in the PCG algorithm (see Algorithm 1) are split across three kernels. This follows the original design of the 3DGS differentiable rasterizer [20]. In both cases, we need to calculate the Jacobian-vector product of the form $\mathbf{g}=\mathbf{J}^{T}\mathbf{u}$ where $\mathbf{J}\in\mathbb{R}^{NxM}$ is the Jacobian matrix of $N$ residuals and $M$ Gaussian parameters and $\mathbf{u}\in\mathbb{R}^{N}$ is an input vector. The $k$-th element in the output vector is calculated as
@@
where $\mathbf{r}_{i}$ is the $i$-th residual and $\mathbf{x}_{k}$ is the $k$-th Gaussian parameter. Following the chain-rule, it is possible to split up the gradient $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}=\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}}$. We can use this to split the computation across three smaller kernels, where only the first needs to calculate the sum over all residuals: $\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\mathbf{u}_{i}$. The other kernels then only calculate the remaining steps by parallelizing over Gaussians. Since the sum needs to be implemented atomically, i.e., multiple threads write to the same output position, this is the main bottleneck for the kernel implementation. By splitting the computation in three smaller parts, we only need to do the costly summation in the first kernel and only to intermediate attributes. Concretely $\mathbf{y}_{k}$ are the 2D mean, color, and opacity attributes of the $k$-th projected Gaussian. Since both the buildCache and applyJT kernels use the gradient cache and our proposed per-pixel-per-splat parallelization pattern, we can implement the sum by first doing a warp reduce and then only issuing one atomicAdd statement per warp.
In contrast, the applyJ and diagJTJ computations cannot be split up into smaller kernels. Concretely, the applyJ kernel calculates $\mathbf{u}=\mathbf{J}\mathbf{p}$ with $\mathbf{p}\in\mathbb{R}^{M}$. The $k$-th element in the output vector is calculated as
@@
Similarly, the diagJTJ kernel calculates $\mathbf{M}=\text{diag}(\mathbf{J}^{T}\mathbf{J})\in\mathbb{R}^{M}$. The $k$-th element in the output vector is calculated as
$\mathbf{g}_{k}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}})^{2}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}})^{2}$ (10)
@@
In both cases it is not possible to move part of the gradients outside of the sum. As a consequence, both the applyJ and diagJTJ are implemented as one kernel, where each thread directly calculates the complete backward path to all Gaussian attributes. This slightly increases the number of required registers and the runtime compared to the applyJT kernel (see Tab. 4).
@@
The applyJ kernel requires the input vector $\mathbf{p}$ to be sorted per Gaussian to make reading from it coalesced. That is: $\mathbf{p}=[x_{1}^{a},...,x_{1}^{z},...,x_{M}^{a},...,x_{M}^{z}]^{T}$, where $x_{k}^{a}$ is the value corresponding to the $a$-th parameter of the $k$-th Gaussian. In total, each Gaussian consists of 59 parameters: 11 for position, rotation, scaling, and opacity and 48 for all Spherical Harmonics coefficients of degree 3. In contrast, all other kernels require the input vector to be sorted per attribute to make reading from it coalesced. That is: $\mathbf{q}=[x_{1}^{a},...,x_{M}^{a},...,x_{1}^{z},...,x_{M}^{z}]^{T}$. We use the structure of $\mathbf{q}$ for all other vector-vector calculations in Algorithm 1 as well. Whenever we call the applyJ kernel, we thus first call the sortX kernel that restructures $\mathbf{q}$ to the layout of $\mathbf{p}$.
Appendix B Derivation of Weights for Subsampling
@@
We sample batches of fewer images to decrease the size of the gradient cache (see Sec. 3.3.2). To combine the update vectors from multiple batches, we calculate the weighted mean, as detailed in Eq. 7. This weighted mean approximates the true solution without any image subsampling. That is, to obtain the update vector, we have to solve the normal equations, as detailed in Eq. 3. When subsampling images, we split the number of total residuals $M$ into smaller chunks. Let’s consider the case of two chunks, labeled as $1$, $2$. The normal equations (without subsampling) can then be written as:
@@
where we drop the additional LM regularization term for clarity and divide the Jacobian and residual vector into separate matrices according to the chunks. The solution to the normal equations is obtained by:
@@
In contrast, when we subsample images, we obtain two separate solutions as:
@@
Appendix C Detailed Analysis of Runtime
Kernel Duration (ms) $\downarrow$ Compute Throughput (%$\uparrow$ Memory Throughput (%$\uparrow$ Register Count $\downarrow$

### v2

@@ 4.4 Limitations
3DGS [23] (L1/SSIM) 0.862 27.23 0.108 1573
3DGS + Ours (L1/SSIM) 0.863 27.29 0.110 1175
@@
Table 2: Ablation of objective. We compare using the L1/SSIM losses against the L2 loss. For both, 3DGS [23] optimized with ADAM and combined with ours, we achieve better results with the L1/SSIM objective. In both cases, our method accelerates the convergence. Results on the GARDEN scene from MipNeRF360 [4].
@@
Table 3: Ablation of batch-size. Selecting fewer images per LM iteration reduces runtime and consumed GPU memory, while only slightly impacting quality. This demonstrates that image subsampling (Sec. 3.3) is compatible with LM in our task. Results obtained after initialization with 3DGS [23] and with $\text{n}_{\text{b}}{=}1$.
[…]
3DGS [23] 10,000 1 1222 29.51
3DGS [23] 50 75 962 29.54
3DGS [23] 130 75 1193 29.68
+ Ours 5 75 951 29.72
DISTWAR [12] 10,000 1 841 29.47
DISTWAR [12] 50 75 681 29.49
DISTWAR [12] 130 75 814 29.58
+ Ours 5 75 672 29.60
gsplat [48] 10,000 1 919 29.52
gsplat [48] 50 75 724 29.53
gsplat [48] 130 75 892 29.56
+ Ours 5 75 716 29.58
Taming-3DGS [32] 10,000 1 447 29.84
Taming-3DGS [32] 50 75 328 29.86
Taming-3DGS [32] 130 75 391 29.91
+ Ours 5 75 347 29.91
Table 4: Analysis of multi-view constraints. We obtain higher quality update steps from our LM optimization and need fewer iterations to converge. Using equally many images in a batch, baselines using ADAM still require more iterations and runtime to reach similar quality. Results averaged on DeepBlending [19].
@@
Figure 6: Runtime Analysis. One iteration of our LM optimizer is dominated by solving PCG and building the cache. Measured on the GARDEN scene from Mip-NeRF360 [4] after densification.
@@
By replacing ADAM with our LM scheme, we accelerate the 3DGS convergence speed by 20% on average for all datasets and baselines. However, some drawbacks remain. First, our approach requires more GPU memory than baselines, due to our gradient cache (Sec. 3.3). Depending on the number and resolution of images, this might require additional CPU offloading of cache parts to run our method on smaller GPUs. Following Mallick et al[32], one can further reduce the cache size by storing the gradients $\frac{\partial c}{\partial s}$ only for every 32nd splat along a ray and re-doing the $\alpha$-blending in these local windows. Second, our two-stage approach relies on ADAM for the densification. 3DGS [23] densifies Gaussians up to 140 times, which is not easily transferable to the granularity of only 5-10 LM iterations. Instead, one could explore and integrate recent alternatives [25, 5, 30].
@@ 5 Conclusion
We have presented 3DGS-LM, a method that accelerates the reconstruction of 3D Gaussian-Splatting [23] by replacing the ADAM optimizer with a tailored Levenberg-Marquardt (LM) (Sec. 3.2). We show that with our data parallelization scheme we can efficiently solve the normal equations with PCG in custom CUDA kernels (Sec. 3.3). Employed in a two-stage approach (Sec. 3.4), this leads to a 20% runtime acceleration compared to baselines. We further demonstrate that our approach is agnostic to other methods [12, 48, 32], which further improves the optimization runtime; i.e., we can easily combine our proposed optimizer with faster 3DGS methods. Overall, we believe that the ability of faster 3DGS reconstructions with our method will open up further research avenues like [28] and make 3DGS more practical across a wide range of real-world applications.
@@ Appendix A More Details About CUDA Kernel Design
We introduce the necessary CUDA kernels to calculate the PCG algorithm in Sec. 3.3. In this section, we provide additional implementation details.
A.1 Parallelization Pattern
We implement the per-pixel-per-splat parallelization pattern in our CUDA kernels by reading subsequent entries from the gradient cache in subsequent threads. This makes reading cache values perfectly coalesced and therefore minimizes the overhead caused by the operation. The gradient cache is sorted over Gaussians, which means that subsequent entries refer to different rays (pixels) that saw this projected Gaussian (splat). In general, one thread handles all residuals corresponding to the respective pixel, i.e., many computations are shared across color channels and are therefore combined.
A.2 Design Of buildCache And applyJT
The necessary computations for the buildCache and applyJT steps in PCG (see Algorithm 1) are split across three kernels. This follows the original design of the 3DGS differentiable rasterizer [23]. In both cases, we calculate the Jacobian-vector product of the form $\mathbf{g}{=}\mathbf{J}^{T}\mathbf{u}$ where $\mathbf{J}{\in}\mathbb{R}^{NxM}$ is the Jacobian matrix of $N$ residuals and $M$ Gaussian parameters and $\mathbf{u}{\in}\mathbb{R}^{N}$ is an input vector. The $k$-th element in the output vector is calculated as:
@@ A.2 Design Of buildCache And applyJT
where $\mathbf{r}_{i}$ is the $i$-th residual and $\mathbf{x}_{k}$ is the $k$-th Gaussian parameter. In other words, $\mathbf{g}_{k}$ is a sum over all residuals for the $k$-th Gaussian parameter. Following the chain-rule, it is possible to split up the gradient $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}{=}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}}$. We can use this to split the computation across three smaller kernels, where only the first needs to calculate the sum over all residuals: $\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\mathbf{u}_{i}$. This sum is the main bottleneck for the kernel implementation, since it needs to be implemented atomically (i.e., multiple threads write to the same output position). The other kernels then only calculate the remaining steps by parallelizing over Gaussians. We slightly abuse notation and denote with $\mathbf{y}_{k}$ the 2D mean, color, and opacity attributes of the $k$-th projected Gaussian. That is, the first kernel sums up the partial derivatives to each of these attributes in separate vectors. In the following, we add the suffixes _p1, _p2, _p3 to denote the three kernels for the respective operation (where _p1 refers to the kernel that calculates the sum over residuals).
The buildCache_p1 utilizes the original per-pixel parallelization, whereas the applyJT_p1 kernel use the gradient cache and our proposed per-pixel-per-splat parallelization pattern. The gradient cache is sorted over Gaussians, i.e., subsequent entries correspond to different rays of the same splat. This allows us to efficiently implement the sum by first performing a segmented warp reduce and then only issuing one atomicAdd statement per warp.
A.3 Design Of applyJ And diagJTJ
In contrast, the applyJ and diagJTJ computations cannot be split up into smaller kernels. Concretely, the applyJ kernel calculates $\mathbf{u}{=}\mathbf{J}\mathbf{p}$ with $\mathbf{p}{\in}\mathbb{R}^{M}$. The $k$-th element in the output vector is calculated as:
@@ A.3 Design Of applyJ And diagJTJ
In other words, $\mathbf{u}_{k}$ is a sum over all Gaussian attributes for the $k$-th residual. Similarly, the diagJTJ kernel calculates $\mathbf{M}=\text{diag}(\mathbf{J}^{T}\mathbf{J})\in\mathbb{R}^{M}$. The $k$-th element in the output vector is calculated as:
$\mathbf{M}_{k}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}})^{2}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}})^{2}$ (11)
@@
In both cases it is not possible to move part of the gradients outside of the sum. As a consequence, both applyJ and diagJTJ are implemented as one kernel, where each thread directly calculates the final partial derivatives to all Gaussian attributes. This slightly increases the number of required registers and the runtime compared to the applyJT kernel (see Tab. 5). The diagJTJ kernel makes use of the same segmented warp reduce as applyJT_p1 for efficiently summing up the squared partial derivatives. The applyJ kernel first sums up over all Gaussian attributes within each thread separately. Then, we only issue one atomicAdd statement for each residual per thread.
@@
The applyJ kernel requires the input vector $\mathbf{p}$ to be sorted per Gaussian to make reading from it coalesced. That is: $\mathbf{p}{=}[x_{1}^{a},...,x_{1}^{z},...,x_{M}^{a},...,x_{M}^{z}]^{T}$, where $x_{k}^{a}$ is the value corresponding to the $a$-th parameter of the $k$-th Gaussian. In total, each Gaussian consists of 59 parameters: 11 for position, rotation, scaling, and opacity and 48 for all Spherical Harmonics coefficients of degree 3. In contrast, all other kernels require the output vector to be sorted per attribute to make writing to it coalesced. That is: $\mathbf{q}{=}[x_{1}^{a},...,x_{M}^{a},...,x_{1}^{z},...,x_{M}^{z}]^{T}$. We use the structure of $\mathbf{q}$ for all other vector-vector calculations in Algorithm 1 as well. Whenever we call the applyJ kernel, we thus first call the sortX kernel that restructures $\mathbf{q}$ to the layout of $\mathbf{p}$.
A.4 Precomputation Of Residual-To-Pixel Weights
We adopt the square root formulation of the residuals in our energy formulation (see Eq. 3). We efficiently precompute the contribution of the square root to the partial derivatives of Eq. 9, Eq. 10, and Eq. 11. In the following, we divide the partial derivatives from the $i$-th residual to the $k$-th Gaussian attribute into two stages:
$\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}=\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ (12)
where $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$ goes from the residual to the rendered pixel color and $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ from the pixel color to the Gaussian attribute. Since we adopt the L1 and SSIM loss terms and take their square root, the terms $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$ need to be calculated accordingly. In contrast, when using the L2 loss they take a trivial form of $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}{=}1$. In the following, we show that we can simplify the calculation of $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}$ in the kernels by precomputing $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$.
The $k$-th element of $\mathbf{g}$ is calculated as:
$\mathbf{g}_{k}=(\mathbf{J}^{T}\mathbf{J}\mathbf{p})_{k}=\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}\sum_{j=0}^{M}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}$ (13)
By substituting Eq. 12 into Eq. 13, we factor out $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$:
$\mathbf{g}_{k}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}})^{2}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}\sum_{j=0}^{M}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}$ (14)
The terms $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ are identical for a residual of the same pixel and color channel that corresponds to either the L1 or SSIM loss terms, respectively. To avoid computing $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ twice (and therefore doubling the grid size of all kernels), we instead sum up the contribution of both loss terms:
$(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}})^{2}=(\frac{\partial\mathbf{r}_{i}^{1}}{\partial\mathbf{c}_{i}})^{2}+(\frac{\partial\mathbf{r}_{i}^{2}}{\partial\mathbf{c}_{i}})^{2}$ (15)
where $\mathbf{r}_{i}^{1}$ corresponds to the $i$-th L1 residual and $\mathbf{r}_{i}^{2}$ to the $i$-th SSIM residual. Additionally, we implement the multiplication with $(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}})^{2}$ in Eq. 14 as elementwise vector product (denoted by $\odot$) of $\mathbf{\hat{u}}{=}[\sum_{j=0}^{M}\frac{\partial\mathbf{c}_{0}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}...\sum_{j=0}^{M}\frac{\partial\mathbf{c}_{N}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}]$ and $\mathbf{\nabla r}{=}[(\frac{\partial\mathbf{r}_{0}}{\partial\mathbf{c}_{0}})^{2}...(\frac{\partial\mathbf{r}_{N}}{\partial\mathbf{c}_{N}})^{2}]$:
$\mathbf{g}_{k}=\sum_{i=0}^{N}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}(\mathbf{\hat{u}\odot\nabla r})_{i}$ (16)
This avoids additional uncoalesced global memory reads to $\mathbf{\nabla r}$ in the CUDA kernels. Instead, we calculate $\mathbf{\hat{u}\odot\nabla r}$ in a separate operation after the applyJ kernel and before applyJT. This also simplifies the kernels, since they now only need to compute $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ instead of $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}$. Therefore, the only runtime overhead of using the L1 and SSIM residual terms over the L2 residuals is the computation of $\mathbf{\nabla r}$. However, this can be efficiently computed using backpropagation (autograd) and is therefore not a bottleneck.
Appendix B Derivation Of Image Subsampling Weights
@@
We subsample batches of images to decrease the size of the gradient cache (see Sec. 3.3). To combine the update vectors from multiple batches, we calculate their weighted mean, as detailed in Eq. 8. This weighted mean approximates the “true” solution to the normal equations (Eq. 4), that does not rely on any image subsampling (and instead uses all available training images). When subsampling images, we split the number of total residuals into smaller chunks. In the following, we consider the case of two chunks (labeled as 1 and 2), but the same applies to any number of chunks. We re-write the normal equations (without subsampling) using the chunk notation as:
@@
where we drop the additional LM regularization term for clarity and divide the Jacobian and residual vector into separate matrices/vectors according to the chunks. The solution to the normal equations is obtained by:
@@
In contrast, when we subsample images, we solve the normal equations separately and obtain two solutions:
@@
Appendix C Detailed Runtime Analysis
[…]